# Towards the Worst-case Robustness of Large Language Models

### v3

@@ B.3 Diffusion Models for Adversarial Robustness
Diffusion models [104, 105] have achieved notable success in defending against visual adversarial examples [106, 107, 108, 109, 110, 111]. In particular, they are widely used as a plug-and-play purification method, named DiffPure, making them suitable for commercial models [112]. As illustrated in Figure 2, given a model to be protected model, $f$, and a diffusion denoiser $D$, DiffPure involves two main steps: First, it adds Gaussian noise with variance $\sigma_{\tau}^{2}$ to the input images, and then denoising these noisy images using the diffusion model $D$.
Intuitively, the norm of the added Gaussian noise is much larger than that of the adversarial perturbations, effectively washing out the adversarial nature of the small-norm perturbations [106]. Theoretically, this procedure not only increases the log-likelihood of input images, pushing them back from out-of-distribution to in-distribution [106, 109], but also implicitly constructs a smooth classifier $g(\bm{x})=\mathbb{E}_{\bm{x}_{\tau}\sim\mathcal{N}(\bm{x},\sigma_{\tau}^{2}\bm{I})}[f(D(\bm{x}_{t}))]$. The mathematical properties of this classifier have been extensively studied, providing theoretical proof on whether adversarial examples can exist within certain neighborhoods [111, 109, 113, 110].
@@ B.4 More Related Work on Certified Robustness
Certified robustness by masking. Certified robustness through masking has been extensively studied in previous work [29, 114, 115, 116] in both text and image domains (e.g., partitioning images into patches and masking them). The certification approach for DiffTextPure-Absorb differs slightly from these works, as tokens are masked with a probability rather than at a fixed ratio, leading to a much more neat result, as shown in Sec. 5.1. Zeng et al. [29] suggest that this certified lower bound can be improved by introducing an auxiliary variable. However, their approach does not incorporate hypothesis testing or account for type-one error in estimating this auxiliary variable. For randomized smoothing via masking, it is obvious that this bound is tight as there exists a worst-case $f$ that fails entirely on region $L_{1}$. When fixing their bound with hypothesis testing using Bonferroni correction, it is clear that this produces the same result.
@@
Certified robustness by random perturbing words. Jia et al. [30] uses interval bounds propagation to propagate the activation bounds to the final layers. These methods currently are not scalable to large models. On the contrary, we adopt randomized smoothing, a model-agnostic certification approach, which is thus more scalable.
@@
Universal certification. Lee et al. [117] also establish a lower bound when smoothing a pre-trained model with randomly perturbed words, but there are several key differences compared to our work. First, we demonstrate that the certified robustness problem can be formulated as a Fractional Knapsack problem, making the approach more intuitive and easier. Second, we show that this can be further improved when the base model $f$ is a hard function, which becomes a 0-1 Knapsack problem and can obtain a stronger result using dynamic programming. What’s more, we greatly simplify the problem by showing that only the different part needs to be considered (see Sec. 5.2), which significantly streamlines the computation of the value-to-weight ratio (see Theorem 5.4). Finally, we show that the uniform kernel reduces to the absorbing kernel when $|\mathcal{V}|\to\infty$, i.e., Figure 1(c) gradually becomes Figure 1(a), giving more theoretical insights.
@@
Certified robustness using synonyms substitution. Ye et al. [28] perturbs words into synonyms (including the original word) with the same probability to achieve certified robustness against word substitution attacks using synonyms. This certified bound closely resembles our DiffTextPure-Absorb method. Specifically, for any perturbed sentence $\bm{z}$, either it cannot result from perturbing the natural or adversarial sentence (trading rate of 0), or it is derived from both with the same probability (trading rate of 1). Consequently, the procedure of certifying using this synonym distribution is the same as that of our absorbing kernel. This approach cannot be generalized to certify word substitution attacks beyond synonyms, as perturbing uniformly into each word in the whole vocabulary with the same probability would completely disrupt the semantics.
@@
Certified robustness for large language models. Kumar et al. [62] first certify large language models against suffix attacks and insertion attacks by randomly deleting tokens. In our notation, they set $p(\bm{z}|\bm{x})$ as a uniform distribution over sentences that have deleted fewer than $k$ tokens from $\bm{x}$, and they set the threshold to infinitesimally small, i.e., as long as there is one harmful $\bm{z}$, they classify $\bm{x}$ as harmful. Therefore, their certified accuracy is exactly the empirical accuracy of detectors on the original text. Since it is extremely easy to achieve 100% TPR on clean data, one will definitely get 100% certified accuracy and +$\infty$+ ∞ certified radius using Kumar et al. [62]. All the randomized smoothing methods degrade to Kumar et al. [62] against suffix attacks when $\beta\to 1$.
@@
Robey et al. [31] propose smoothing a language model by randomly perturbing each character, rather than tokens. They also do not certify their defense. Their theorem is based on an assumption they define themselves, called k-stable, which states that perturbing $k+1$ characters would result in a change. This assumption indeed already implicitly implies robustness. In this work, we do not make any such assumptions. Instead, we certify each input $\bm{x}$ independently, rather than relying on a distribution.
@@ B.5 On Discrete Diffusion Models
Discrete diffusion models extend traditional diffusion models to the discrete domain, enabling the modeling of language inputs [54, 118, 20]. Given a vocabulary $\mathcal{V}=\{1,\cdots,|V|\}$, sequence length $N$, a data distribution $p:=p_{0}\in\mathcal{V}^{N}$, the forward process creates a sequence of distributions $p_{t}$ by randomly perturbing each word according to a continuous-time Markov chain:
@@
Typically, $Q_{t}$ is defined as $\sigma(t)Q$ for simplicity, where $\sigma(t)$ is a monotonic noise schedule designed to ensure that $p_{T}$ approaches a simple prior distribution $p_{prior}$. Eq. 8 provides two frequency choices of $Q$. when $Q=Q^{\rm uniform}$, this Markov chain progressively and uniformly perturbs each word to any other word over time. When $Q=Q^{\rm absorb}$, it gradually perturbs each word into an absorbing token.
@@
The forward process has an analytical form due to its simplicity. For the $i$-th word $\bm{x}_{0}^{i}$, $p_{t|0}(\cdot|\bm{x}_{0}^{i})=\exp(\int_{0}^{t}\sigma(s)dsQ)_{\bm{x}_{0}^{i}}$. It also has a well-known reversal given by another diffusion matrix $\overline{Q}_{t}$ [119]. For the $i$-th word, the reversal is:
[…]
where $\bm{y}$ is another sentence that differs from $\bm{x}_{t}$ only at $i$-th position, $\frac{p_{t}(\bm{y})}{p_{t}(\bm{x})}$ is referred as the concrete score. Once we train a score network $s_{\theta}(\bm{x}_{t},t)$ to approximate the concrete score, we can sample new instances using Eq. 9 by substituting the unknown score $\frac{p_{t}(\bm{y})}{p_{t}(\bm{x})}$ with the neural network-estimated score $s_{\theta}(\bm{x},t)$ [54, 20]. Unlike the forward process, the reverse process lacks an analytical form due to the involvement of a neural network. Consequently, numerical methods such as an Euler solver or a $\tau$-leaping solver are typically employed to approximate the backward Markov chain.
@@ B.6 Certification on Different Tasks
Case 1: Image Classification. In image classification, $f$ can be a classifier mapping from the image domain to one interested class in $K-1$ probability simplex. The smoothing distribution $p(\bm{z}|\bm{x})$ can be a Gaussian distribution [23, 113], a Uniform distribution [120, 117], Laplacian distribution [27], or other types of distributions. If we can certify that $p_{adv}\geq 0.5$ in Definition 4.1, it guarantees that the classifier will consistently produce the correct result for all $\bm{x}_{adv}$ satisfying $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$. This is because the probability of the true class remains the highest among all output probabilities.
Case 2: Text Classification. Similarly, for a text classifier $f:\mathcal{V}^{N}\to[0,1]$, that maps from a text to a probability of outputting a target class that we are interested in, the smoothing distribution can be derived from the noisy process of diffusion models, $p_{t|0}(\bm{x}_{\tau}|\bm{x})$, such as randomly replacing or masking words [20], as described in Sec. F.2. If we can certify that $p_{adv}\geq 0.5$ for the correct class $y$, it ensures that $y$ remains the largest output of $g(\bm{x}_{adv})$, guaranteeing robust classification.
Case 3: Text Safety. This has already been extensively discussed in Sec. 4.1.
[…]
@@ Appendix C Proofs for Knapsack Solvers
C.1 Proof of Theorem 4.3
The optimality of the greedy algorithm in Theorem 4.3 has been extensively proven [50, 51]. The proof is typically conducted by contradiction. By sorting the items by their value-to-weight ratio, assume that there exists a better selection than the one obtained by selecting items based on their value-to-weight ratios. Comparing the differing items in these two selections, both must have the same volume, but the selection based on value-to-weight ratio will always have a higher ratio, and thus a higher value. Therefore, in the fractional knapsack problem, it is impossible to find a better approach than selecting items in descending order of their value-to-weight ratio.
Another proof, more closely related to the approach in [23], uses the method of Lagrange multipliers. Our goal is to find the minimal solution to a constrained optimization problem:
@@
The hypothesis set of the base function $f$ consists of all bounded functions. Normalizing them to $[0,1]$, we define the hypothesis set as =$\mathcal{F}=\{f:\mathcal{X}\to[0,1]\}$. Thus, each $f(\bm{z})$ can take any value in $[0,1]$. We treat $f(\bm{z})$ for each $\bm{z}$ as a variable and compute the derivative of $\mathcal{L}$ with respect to $f(\bm{z})$. For each $\bm{z}$, we have (total of $|\mathcal{X}|$):
@@
In other words, if the value-to-weight ratio $-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ is less than $\lambda$, then $f(\bm{z})$ should be set to 0. If the value-to-weight ratio $-\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ is greater than $\lambda$, then $f(\bm{z})$ should be set to 1. Therefore, the algorithm to solve this problem is to first sort the value-to-weight ratios and then set the corresponding function values to 1 in order, until the constraint $g(\bm{x})=p_{A}$ is satisfied (which controls $\lambda$).
@@
Further narrowing of the hypothesis set can yield better solutions for this constrained optimization problem, e.g., restricting to binary functions =$\mathcal{F}=\{f:\mathcal{X}\to\{0,1\}\}$ or functions with Lipschitz continuity [52, 53].
@@
In Sec. 4.2, we mentioned the connection between the randomized smoothing problem and the 0-1 Knapsack problem. Specifically, if we restrict the hypothesis set of the function $f$ to hard functions that only output binary values (i.e., functions that map to $\{0,1\}$), then the problem at hand becomes a 0-1 Knapsack problem. This restriction leads to a more efficient solution where we can apply dynamic programming to obtain a tighter bound on the robustness of the function.
@@
Given a probability distribution $p(\bm{z}|\bm{x})$ that represents the weight (or quality) of each item, and a corresponding adversarial distribution $-p(\bm{z}|\bm{x}_{adv})$ that represents the value (or profit) of each item, we are tasked with selecting a subset of items such that the total weight (i.e., the total probability mass at the clean example) does not exceed a given threshold $p_{A}$. The goal is to maximize the total value, which is the sum of the negative log-probabilities from the adversarial distribution.
This scenario naturally translates into the 0-1 Knapsack problem, where weights are given by $p(\bm{z}|\bm{x})$, values are given by $-p(\bm{z}|\bm{x}_{adv})$, the capacity of the knapsack is $p_{A}$, and the objective is to maximize the total value, subject to the constraint on the total weight.
@@
The dynamic programming solution is demonstrated in Algorithm 2. It first define $dp[i][w]$ to be the maximum value that can be obtained by considering the first $i$ items, with a knapsack capacity of $w$. For each item $\bm{z}_{i}$, if we can add it to the knapsack (i.e., if the current weight $w$ is greater than or equal to the weight of the item $p(\bm{z}_{i}|\bm{x})$), we update the DP table by considering both the inclusion and exclusion of the item:
@@
Finally, we check whether the maximum value obtained is greater than or equal to the threshold $\tau$. If $-V_{\text{max}}\geq\tau$, then we can certify that the function is provably robust for all distributions with $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$. Otherwise, the function does not meet the robustness criterion.

### v4

@@ B.3 Diffusion Models for Adversarial Robustness
Diffusion models Song et al. (2021); Dhariwal & Nichol (2021) have achieved notable success in defending against visual adversarial examples Nie et al. (2022); Wang et al. (2022); Li et al. (2024b); Xiao et al. (2023); Zhang et al. (2023); Carlini et al. (2023b). In particular, they are widely used as a plug-and-play purification method, named DiffPure, making them suitable for commercial models Zhang et al. (2024a). As illustrated in Figure˜2, given a model to be protected model, $f$, and a diffusion denoiser $D$, DiffPure involves two main steps: First, it adds Gaussian noise with variance $\sigma_{\tau}^{2}$ to the input images, and then denoising these noisy images using the diffusion model $D$.
Intuitively, the norm of the added Gaussian noise is much larger than that of the adversarial perturbations, effectively washing out the adversarial nature of the small-norm perturbations Nie et al. (2022). Theoretically, this procedure not only increases the log-likelihood of input images, pushing them back from out-of-distribution to in-distribution Nie et al. (2022); Xiao et al. (2023), but also implicitly constructs a smooth classifier $g(\bm{x})=\mathbb{E}_{\bm{x}_{\tau}\sim\mathcal{N}(\bm{x},\sigma_{\tau}^{2}\bm{I})}[f(D(\bm{x}_{t}))]$. The mathematical properties of this classifier have been extensively studied, providing theoretical proof on whether adversarial examples can exist within certain neighborhoods Carlini et al. (2023b); Xiao et al. (2023); Chen et al. (2024b); Zhang et al. (2023).
@@ B.4 More Related Work on Certified Robustness
Certified robustness by masking. Certified robustness through masking has been extensively studied in previous work Zeng et al. (2023); Levine & Feizi (2020); Moon et al. (2023); Zhang et al. (2019) in both text and image domains (e.g., partitioning images into patches and masking them). The certification approach for DiffTextPure-Absorb differs slightly from these works, as tokens are masked with a probability rather than at a fixed ratio, leading to a much more neat result, as shown in Sec.˜5.1. Zeng et al. (2023) suggest that this certified lower bound can be improved by introducing an auxiliary variable. However, their approach does not incorporate hypothesis testing or account for type-one error in estimating this auxiliary variable. For randomized smoothing via masking, it is obvious that this bound is tight as there exists a worst-case $f$ that fails entirely on region $L_{1}$. When fixing their bound with hypothesis testing using Bonferroni correction, it is clear that this produces the same result.
@@
Certified robustness by random perturbing words. Jia et al. (2019) uses interval bounds propagation to propagate the activation bounds to the final layers. These methods currently are not scalable to large models. On the contrary, we adopt randomized smoothing, a model-agnostic certification approach, which is thus more scalable.
@@
Universal certification. Lee et al. (2019) also establish a lower bound when smoothing a pre-trained model with randomly perturbed words, but there are several key differences compared to our work. First, we demonstrate that the certified robustness problem can be formulated as a Fractional Knapsack problem, making the approach more intuitive and easier. Second, we show that this can be further improved when the base model $f$ is a hard function, which becomes a 0-1 Knapsack problem and can obtain a stronger result using dynamic programming. What’s more, we greatly simplify the problem by showing that only the different part needs to be considered (see Sec.˜5.2), which significantly streamlines the computation of the value-to-weight ratio (see Theorem˜5.4). Finally, we show that the uniform kernel reduces to the absorbing kernel when $|\mathcal{V}|\to\infty$, i.e., Figure˜1(c) gradually becomes Figure˜1(a), giving more theoretical insights.
@@
Certified robustness using synonyms substitution. Ye et al. (2020) perturbs words into synonyms (including the original word) with the same probability to achieve certified robustness against word substitution attacks using synonyms. This certified bound closely resembles our DiffTextPure-Absorb method. Specifically, for any perturbed sentence $\bm{z}$, either it cannot result from perturbing the natural or adversarial sentence (trading rate of 0), or it is derived from both with the same probability (trading rate of 1). Consequently, the procedure of certifying using this synonym distribution is the same as that of our absorbing kernel. This approach cannot be generalized to certify word substitution attacks beyond synonyms, as perturbing uniformly into each word in the whole vocabulary with the same probability would completely disrupt the semantics.
@@
Certified robustness for large language models. Kumar et al. (2023) first certify large language models against suffix attacks and insertion attacks by randomly deleting tokens. In our notation, they set $p(\bm{z}|\bm{x})$ as a uniform distribution over sentences that have deleted fewer than $k$ tokens from $\bm{x}$, and they set the threshold to infinitesimally small, i.e., as long as there is one harmful $\bm{z}$, they classify $\bm{x}$ as harmful. Therefore, their certified accuracy is exactly the empirical accuracy of detectors on the original text. Since it is extremely easy to achieve 100% TPR on clean data, one will definitely get 100% certified accuracy and $\infty$ certified radius using Kumar et al. (2023). All the randomized smoothing methods degrade to Kumar et al. (2023) against suffix attacks when $\beta\to 1$.
@@
Robey et al. (2023) propose smoothing a language model by randomly perturbing each character, rather than tokens. They also do not certify their defense. Their theorem is based on an assumption they define themselves, called k-stable, which states that perturbing $k+1$ characters would result in a change. This assumption indeed already implicitly implies robustness. In this work, we do not make any such assumptions. Instead, we certify each input $\bm{x}$ independently, rather than relying on a distribution.
@@ B.5 On Discrete Diffusion Models
Discrete diffusion models extend traditional diffusion models to the discrete domain, enabling the modeling of language inputs Meng et al. (2022); Campbell et al. (2022); Lou et al. (2023). Given a vocabulary $\mathcal{V}=\{1,\cdots,|V|\}$, sequence length $N$, a data distribution $p:=p_{0}\in\mathcal{V}^{N}$, the forward process creates a sequence of distributions $p_{t}$ by randomly perturbing each word according to a continuous-time Markov chain:
@@
Typically, $Q_{t}$ is defined as $\sigma(t)Q$ for simplicity, where $\sigma(t)$ is a monotonic noise schedule designed to ensure that $p_{T}$ approaches a simple prior distribution $p_{prior}$. Eq.˜9 provides two frequency choices of $Q$. when $Q=Q^{\rm uniform}$, this Markov chain progressively and uniformly perturbs each word to any other word over time. When $Q=Q^{\rm absorb}$, it gradually perturbs each word into an absorbing token.
@@
The forward process has an analytical form due to its simplicity. For the $i$-th word $\bm{x}_{0}^{i}$, $p_{t|0}(\cdot|\bm{x}_{0}^{i})=\exp(\int_{0}^{t}\sigma(s)dsQ)_{\bm{x}_{0}^{i}}$. It also has a well-known reversal given by another diffusion matrix $\overline{Q}_{t}$ Kelly (2011). For the $i$-th word, the reversal is:
$\frac{dp_{T-t}}{dt}=\overline{Q}_{T-t}p_{T-t},$ $\text{ where }\;\overline{Q}_{t}(\bm{y}^{i},\bm{x}_{t}^{i})=\frac{p_{t}(\bm{y})}{p_{t}(\bm{x})}Q_{t}(\bm{x}_{t}^{i},\bm{y}^{i})\;$ (10)
$\text{ and }\;\overline{Q}_{t}(\bm{x}_{t}^{i},\bm{x}_{t}^{i})=-\sum_{\bm{y}\neq\bm{x}}\overline{Q}_{t}(\bm{y}^{i},\bm{x}_{t}^{i}),$
where $\bm{y}$ is another sentence that differs from $\bm{x}_{t}$ only at $i$-th position, $\frac{p_{t}(\bm{y})}{p_{t}(\bm{x})}$ is referred as the concrete score. Once we train a score network $s_{\theta}(\bm{x}_{t},t)$ to approximate the concrete score, we can sample new instances using Eq.˜10 by substituting the unknown score $\frac{p_{t}(\bm{y})}{p_{t}(\bm{x})}$ with the neural network-estimated score $s_{\theta}(\bm{x},t)$ Meng et al. (2022); Lou et al. (2023). Unlike the forward process, the reverse process lacks an analytical form due to the involvement of a neural network. Consequently, numerical methods such as an Euler solver or a $\tau$-leaping solver are typically employed to approximate the backward Markov chain.
@@ B.6 Certification on Different Tasks
Case 1: Image Classification. In image classification, $f$ can be a classifier mapping from the image domain to one interested class in K1K-1 probability simplex. The smoothing distribution $p(\bm{z}|\bm{x})$ can be a Gaussian distribution Cohen et al. (2019); Chen et al. (2024b), a Uniform distribution Levine & Feizi (2021); Lee et al. (2019), Laplacian distribution Teng et al. (2020), or other types of distributions. If we can certify that $p_{adv}\geq 0.5$ in Eq.˜1, it guarantees that the classifier will consistently produce the correct result for all $\bm{x}_{adv}$ satisfying $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$. This is because the probability of the true class remains the highest among all output probabilities.
Case 2: Multi-class classification [Huanran: TODO]
Case 2: Text Classification. Similarly, for a text classifier $f:\mathcal{V}^{N}\to[0,1]$, that maps from a text to a probability of outputting a target class that we are interested in, the smoothing distribution can be derived from the noisy process of diffusion models, $p_{t|0}(\bm{x}_{\tau}|\bm{x})$, such as randomly replacing or masking words Lou et al. (2023), as described in Sec.˜F.2. If we can certify that $p_{adv}\geq 0.5$ for the correct class $y$, it ensures that $y$ remains the largest output of $g(\bm{x}_{adv})$, guaranteeing robust classification.
Case 3: Text Safety. This has already been extensively discussed in Sec.˜4.1.
Case 4: DiffTextPure. Given a bounded base function $\hat{f}:\mathcal{X}\to[0,1]$, DiffTextPure set $f:=\hat{f}\circ D$, where $D$ is the denoiser, and construct the smoothed function $g(\bm{x})=\mathbb{E}_{p(\bm{z|\bm{x}})}[f(\bm{z})]$. Therefore, DiffTextPure do not require fine-tuning base model $\hat{f}$ on noisy distribution $p(\bm{z})=\int p(\bm{z|\bm{x}})p(\bm{x})d\bm{x}$.
@@ Appendix C Proofs for Knapsack Solvers
C.1 Proof of Theorem˜4.3
The optimality of the greedy algorithm in Theorem˜4.3 has been extensively proven Aho & Hopcroft (1974); Cormen et al. (2022). The proof is typically conducted by contradiction. By sorting the items by their value-to-weight ratio, assume that there exists a better selection than the one obtained by selecting items based on their value-to-weight ratios. Comparing the differing items in these two selections, both must have the same volume, but the selection based on value-to-weight ratio will always have a higher ratio, and thus a higher value. Therefore, in the fractional knapsack problem, it is impossible to find a better approach than selecting items in descending order of their value-to-weight ratio.
Another proof, more closely related to the approach in Cohen et al. (2019), uses the method of Lagrange multipliers. Our goal is to find the minimal solution to a constrained optimization problem:
@@
The hypothesis set of the base function $f$ consists of all bounded functions. Normalizing them to $[0,1]$, we define the hypothesis set as $\mathcal{F}=\{f:\mathcal{X}\to[0,1]\}$. Thus, each $f(\bm{z})$ can take any value in $[0,1]$. We treat $f(\bm{z})$ for each $\bm{z}$ as a variable and compute the derivative of $\mathcal{L}$ with respect to $f(\bm{z})$. For each $\bm{z}$, we have (total of $|\mathcal{X}|$):
@@
In other words, if the value-to-weight ratio $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ is less than $\lambda$, then $f(\bm{z})$ should be set to 0. If the value-to-weight ratio $\frac{p(\bm{z}|\bm{x}_{adv})}{p(\bm{z}|\bm{x})}$ is greater than $\lambda$, then $f(\bm{z})$ should be set to 1. Therefore, the algorithm to solve this problem is to first sort the value-to-weight ratios and then set the corresponding function values to 1 in order, until the constraint $g(\bm{x})=p_{A}$ is satisfied (which controls $\lambda$).
@@
Further narrowing of the hypothesis set can yield better solutions for this constrained optimization problem, e.g., restricting to binary functions $\mathcal{F}=\{f:\mathcal{X}\to\{0,1\}\}$ or functions with Lipschitz continuity Chen et al. (2024a); Delattre et al. (2024).
@@
In Sec.˜4.2, we mentioned the connection between the randomized smoothing problem and the 0-1 Knapsack problem. Specifically, if we restrict the hypothesis set of the function $f$ to hard functions that only output binary values (i.e., functions that map to $\{0,1\}$), then the problem at hand becomes a 0-1 Knapsack problem. This restriction leads to a more efficient solution where we can apply dynamic programming to obtain a tighter bound on the robustness of the function.
@@
Given a probability distribution $p(\bm{z}|\bm{x})$ that represents the weight (or quality) of each item, and a corresponding adversarial distribution $p(\bm{z}|\bm{x}_{adv})$ that represents the value (or profit) of each item, we are tasked with selecting a subset of items such that the total weight (i.e., the total probability mass at the clean example) does not exceed a given threshold $p_{A}$. The goal is to maximize the total value, which is the sum of the negative log-probabilities from the adversarial distribution.
This scenario naturally translates into the 0-1 Knapsack problem, where weights are given by $p(\bm{z}|\bm{x})$, values are given by $p(\bm{z}|\bm{x}_{adv})$, the capacity of the knapsack is $p_{A}$, and the objective is to maximize the total value, subject to the constraint on the total weight.
@@
The dynamic programming solution is demonstrated in Algorithm˜2. It first define $dp[i][w]$ to be the maximum value that can be obtained by considering the first $i$ items, with a knapsack capacity of $w$. For each item $\bm{z}_{i}$, if we can add it to the knapsack (i.e., if the current weight $w$ is greater than or equal to the weight of the item $p(\bm{z}_{i}|\bm{x})$), we update the DP table by considering both the inclusion and exclusion of the item:
@@
Finally, we check whether the maximum value obtained is greater than or equal to the threshold $\tau$. If $V_{\text{max}}\geq\tau$, then we can certify that the function is provably robust for all distributions with $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$. Otherwise, the function does not meet the robustness criterion.